Main INDEX
Monthly INDEX
PREV
NEXT
User name A. Puckett
Log entry time 02:20:34 on April25,2007
Entry number 199947
Followups:
keyword=Left shield hut temp. looks stable.
The temperature appears to have stabilized for now at 83.9 deg. Heidi did not answer my page, but the most recent detector checks look fine, all expected channels have good data and for the past hour or so the temp. reading hasn't changed. So unless something changes I won't bother the techs at this hour.